OpenStudy (anonymous):

how do you solve -3x^1/3+4x^1/2+5x^1/3+4x^1/2 ?

OpenStudy (anonymous):

\[\frac{ -3x }{ 3+4x^{1/2} }*\frac{ 3+4x^{1/2} }{ 2+5x }=\frac{ -3x }{ 2+5x }\]